Someone solve these two problems for me plz: ( w/ explanantions plz) or answers if u dont have time

1)A sugar solution was made by mixing 9 L of a 30% sugar solution and 1 L of a 20% sugar solution. What is the concentration of the sugar mixture?

2) 1 yd of soil containing 40 % sand was mixed into 4 yd of soil containing 20% sand. what is the sand content of the mixture?

Answer by stanbon(75887)

1)
A sugar solution was made by mixing 9 L of a 30% sugar solution and 1 L of a 20% sugar solution. What is the concentration of the sugar mixture?
----
sugar + sugar = sugar
0.30*9 + 0.20*1 = x(9+1)
-----
Multiply thru by 100 to get:
30*9 + 20*1 = x(1000)
1000x = 270+20
1000x = 290
x = 0.29 = 29%
===============================
2) 1 yd of soil containing 40 % sand was mixed into 4 yd of soil containing 20% sand. what is the sand content of the mixture?
sand + sand = sans
0.40*1 + 0.20*4 = x(1+4)
40*1 + 20*4 = 500x
500x = 120
x = 0.24 = 24% of the 5 yds of soil is sand
============
Cheers,
Stan H.
=============